Solve these question on graphical representation of linear equations including the graphs for each question
1. Plot the graph of the equations 2x - 3y = 6 and 2x + y = -10 and interpret the result.
2. Plot the graph of the equations 2x + 4y = 10 and 3x + 6y = 12 and interpret the result.
3. Determine graphically the vertices of the triangle, the equation of whose sides are given as y = x; y = 0; 2x + 3y = 10. Interpret the result

Answer by CrazyMan Jr.(21)

Here's the answer to question #1
2x-3y = 6 -4y = 16
2x+y = -10 -4y/-4 = 16/-4 y = -4
Sub In
2x-3(-4) = 6
2x+12 = 6
2x = -6
2x/2 = -6/2
x = -3
Sub In Again
2(-3)+(-4) = -10
-6+-4 = -10
-10 = -10
J-Man